# OPTIMA: Optimal One-Shot Pruning for LLMs via Quadratic Programming Reconstruction

### v1

@@ Abstract
Post-training model pruning is a promising solution, yet it faces a trade-off: simple heuristics that zero weights are fast but degrade accuracy, while principled joint optimization methods recover accuracy but are computationally infeasible at modern scale. One-shot methods such as SparseGPT offer a practical trade-off in optimality by applying efficient, approximate heuristic weight updates. To close this gap, we introduce OPTIMA, a practical one-shot post-training pruning method that balances accuracy and scalability. OPTIMA casts layer-wise weight reconstruction after mask selection as independent, row-wise Quadratic Programs (QPs) that share a common layer Hessian. Solving these QPs yields the per-row globally optimal update with respect to the reconstruction objective given the estimated Hessian. The shared-Hessian structure makes the problem highly amenable to batching on accelerators. We implement an accelerator-friendly QP solver that accumulates one Hessian per layer and solves many small QPs in parallel, enabling one-shot post-training pruning at scale on a single accelerator without fine-tuning. OPTIMA integrates with existing mask selectors and consistently improves zero-shot performance across multiple LLM families and sparsity regimes, yielding up to 3.97% absolute accuracy improvement. On an NVIDIA H100, OPTIMA prunes a 8B-parameter transformer end-to-end in 40 hours with 60 GB peak memory. Together, these results set a new state-of-the-art accuracy-efficiency trade-offs for one-shot post-training pruning. [^1] The code and data for OPTIMA is available at https://github.com/paramathic/optima
@@ 1 Introduction
In this paper we introduce OPTIMA, a practical one-shot post-training pruning framework that closes much of this gap by combining principled optimality with accelerator-grade efficiency. The core idea is a precise reformulation of the layer-wise reconstruction step that follows mask selection. That is, after fixing a binary mask for a weight matrix, the reconstruction (least-squares) objective decomposes across rows and each row’s update can be written as a small quadratic program (QP). Crucially, every row in the same layer shares the same Hessian matrix $H=X^{\top}X$, while the linear constraints differ only according to which entries the mask removes. This shared-Hessian, row-wise QP structure yields two immediate benefits: (1) per-row global optimality for the reconstruction objective (given the estimated Hessian), and (2) uniform problem structure that enables massive batching and parallelism on off-the-shelf ML accelerators (GPUs/TPUs).
Realizing this formulation in practice requires careful numerical and systems engineering. We adopt a first-order primal–dual QP solver (rAPDHG (pdqp)) that is well-suited to our constrained problems and whose critical operations reduce to matrix–vector products with the shared Hessian. This makes the inner loops extremely efficient on accelerators. We further avoid explicit dense equality matrices by enforcing fixed entries via tight bounds, accumulate layer Hessians incrementally from calibration sequences to save memory, and solve rows in batches so thousands of small QPs are processed in parallel. These implementation choices make OPTIMA not only theoretically principled but also practical to run on a single accelerator.
@@
  • We present a row-wise QP reformulation of the post-training reconstruction problem that yields per-row global optimality under a shared-Hessian model and is provably equivalent to the least-squares objective after mask selection (section 3).
  • We design and implement an accelerator-friendly QP solver pipeline that accumulates a single Hessian per layer, enforces mask constraints via bounds, batches thousands of row QPs, and leverages rAPDHG/MPAX for efficient execution on GPUs/TPUs (detailed in Algorithm 1).
@@
Figure 1: OPTIMA generates a shared Hessian among the different rows of the pruned weight using a small calibration dataset. Then, the weights in different rows will be updated in parallel using a QP solver and the shared Hessian.
@@ 3.4 Efficient implementation
Algorithm 1 Layer-wise Pruning with Batched Row-wise Quadratic Programming
@@
12   Partition the rows of $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{W}$ into batches of size $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}B$.
13   for each batch of rows $\{\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{w}_{j}\}_{j=1}^{B}$ in parallel do
14     for each row $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{w}_{j}$ in the batch do
@@ 5 Conclusion
OPTIMA reformulates post-training weight reconstruction as batched, row-wise Quadratic Programs (QPs) that share a layer Hessian. This yields per-row optimal updates for the reconstruction (least-squares) objective given the estimated Hessian, and the shared-Hessian structure enables massive GPU/TPU parallelism. We implement OPTIMA using an accelerator-friendly primal–dual solver and batched solves of many small per-row QPs (i.e., parallel per-row optimization). OPTIMA functions as a practical, drop-in weight-update step for common mask selectors (Wanda, SparseGPT, Thanos). In our experiments on a single NVIDIA H100, OPTIMA improves zero-shot accuracy across LLM families by up to 3.97% points. These gains hold at high sparsity levels ($\geq 60\%$) and require no post-pruning fine-tuning. Together, these results deliver a principled and scalable approach to accurate one-shot post-training pruning.
@@ Appendix B Comparison with alternative optimizers
Table 7 illustrates this on Gemma 3 1B and OPT 125M (opt) under 50% unstructured sparsity. We show two examples in Table 7, showing that ADAM results to suboptimal solutions. To further test the limitations of optimizers without convergence guarantees, we test ADAM on OPT-125M, and observe that it leads to divergence of the model. On Gemma 3 1B, ADAM yields competitive results in some cases (e.g., slightly lower perplexity for SparseGPT+ADAM at 27.12 versus OPTIMA’s 27.35), but OPTIMA achieves higher overall accuracy (e.g., 44.01% for Wanda+OPTIMA versus 43.72% for Wanda+ADAM). However, on smaller models like OPT 125M, ADAM exhibits instability, leading to divergence and dramatically higher perplexity (e.g., 205.82 for Wanda+ADAM versus 35.44 for Wanda+OPTIMA). This underscores the risks of using non-specialized optimizers for our row-wise QPs, where suboptimal or unstable solutions can degrade model quality. OPTIMA’s use of provably convergent methods like rAPDHG ensures reliable and superior weight updates, making it a more robust choice for post-training pruning.

### v2

@@ Abstract
Post-training model pruning is a promising solution, yet it faces a trade-off: simple heuristics that zero weights are fast but degrade accuracy, while principled joint optimization methods recover accuracy but are computationally infeasible at modern scale. One-shot methods such as SparseGPT offer a practical trade-off in optimality by applying efficient, approximate heuristic weight updates. To close this gap, we introduce OPTIMA, a practical one-shot post-training pruning method that balances accuracy and scalability. OPTIMA casts layer-wise weight reconstruction after mask selection as independent, column-wise Quadratic Programs (QPs) that share a common layer Hessian. Solving these QPs yields the per-column globally optimal update with respect to the reconstruction objective given the estimated Hessian. The shared-Hessian structure makes the problem highly amenable to batching on accelerators. We implement an accelerator-friendly QP solver that accumulates one Hessian per layer and solves many small QPs in parallel, enabling one-shot post-training pruning at scale on a single accelerator without fine-tuning. OPTIMA integrates with existing mask selectors and consistently improves zero-shot performance across multiple LLM families and sparsity regimes, yielding up to 3.97% absolute accuracy improvement. On an NVIDIA H100, OPTIMA prunes a 8B-parameter transformer end-to-end in 40 hours with 60 GB peak memory. Together, these results set a new state-of-the-art accuracy-efficiency trade-offs for one-shot post-training pruning. [^1] The code and data for OPTIMA is available at https://github.com/paramathic/optima
@@ 1 Introduction
In this paper we introduce OPTIMA, a practical one-shot post-training pruning framework that closes much of this gap by combining principled optimality with accelerator-grade efficiency. The core idea is a precise reformulation of the layer-wise reconstruction step that follows mask selection. That is, after fixing a binary mask for a weight matrix, the reconstruction (least-squares) objective decomposes across columns and each column’s update can be written as a small quadratic program (QP). Crucially, every column in the same layer shares the same Hessian matrix $H=X^{\top}X$, while the linear constraints differ only according to which entries the mask removes. This shared-Hessian, column-wise QP structure yields two immediate benefits: (1) per-column global optimality for the reconstruction objective (given the estimated Hessian), and (2) uniform problem structure that enables massive batching and parallelism on off-the-shelf ML accelerators (GPUs/TPUs).
Realizing this formulation in practice requires careful numerical and systems engineering. We adopt a first-order primal–dual QP solver (rAPDHG (pdqp)) that is well-suited to our constrained problems and whose critical operations reduce to matrix–vector products with the shared Hessian. This makes the inner loops extremely efficient on accelerators. We further avoid explicit dense equality matrices by enforcing fixed entries via tight bounds, accumulate layer Hessians incrementally from calibration sequences to save memory, and solve columns in batches so thousands of small QPs are processed in parallel. These implementation choices make OPTIMA not only theoretically principled but also practical to run on a single accelerator.
@@
  • We present a column-wise QP reformulation of the post-training reconstruction problem that yields per-column global optimality under a shared-Hessian model and is provably equivalent to the least-squares objective after mask selection (section 3).
  • We design and implement an accelerator-friendly QP solver pipeline that accumulates a single Hessian per layer, enforces mask constraints via bounds, batches thousands of column QPs, and leverages rAPDHG/MPAX for efficient execution on GPUs/TPUs (detailed in Algorithm 1).
@@
Figure 1: OPTIMA generates a shared Hessian among the different columns of the pruned weight using a small calibration dataset. Then, the weights in different columns will be updated in parallel using a QP solver and the shared Hessian.
@@ 3.4 Efficient implementation
Algorithm 1 Layer-wise Pruning with Batched Column-wise Quadratic Programming
@@
12   Partition the columns of $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{W}$ into batches of size $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}B$.
13   for each batch of columns $\{\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{w}_{j}\}_{j=1}^{B}$ in parallel do
14     for each column $\color[rgb]{0.5,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0.5}\mathbf{w}_{j}$ in the batch do
@@ 5 Conclusion
OPTIMA reformulates post-training weight reconstruction as batched, column-wise Quadratic Programs (QPs) that share a layer Hessian. This yields per-column optimal updates for the reconstruction (least-squares) objective given the estimated Hessian, and the shared-Hessian structure enables massive GPU/TPU parallelism. We implement OPTIMA using an accelerator-friendly primal–dual solver and batched solves of many small per-column QPs (i.e., parallel per-column optimization). OPTIMA functions as a practical, drop-in weight-update step for common mask selectors (Wanda, SparseGPT, Thanos). In our experiments on a single NVIDIA H100, OPTIMA improves zero-shot accuracy across LLM families by up to 3.97% points. These gains hold at high sparsity levels ($\geq 60\%$) and require no post-pruning fine-tuning. Together, these results deliver a principled and scalable approach to accurate one-shot post-training pruning.
@@ Appendix B Comparison with alternative optimizers
Table 7 illustrates this on Gemma 3 1B and OPT 125M (opt) under 50% unstructured sparsity. We show two examples in Table 7, showing that ADAM results to suboptimal solutions. To further test the limitations of optimizers without convergence guarantees, we test ADAM on OPT-125M, and observe that it leads to divergence of the model. On Gemma 3 1B, ADAM yields competitive results in some cases (e.g., slightly lower perplexity for SparseGPT+ADAM at 27.12 versus OPTIMA’s 27.35), but OPTIMA achieves higher overall accuracy (e.g., 44.01% for Wanda+OPTIMA versus 43.72% for Wanda+ADAM). However, on smaller models like OPT 125M, ADAM exhibits instability, leading to divergence and dramatically higher perplexity (e.g., 205.82 for Wanda+ADAM versus 35.44 for Wanda+OPTIMA). This underscores the risks of using non-specialized optimizers for our column-wise QPs, where suboptimal or unstable solutions can degrade model quality. OPTIMA’s use of provably convergent methods like rAPDHG ensures reliable and superior weight updates, making it a more robust choice for post-training pruning.